i have to use distance formula to find the distance between point a which is
(-2,5) and point b which is (4,-3).
how do i do that?

Answer by checkley71(8403)

THE BEST SET OF TOOLS FOR THIS TYPE OF PROBLEM IS A PENCIL/PEN, GRAPH PAPER & A STRAIGHT EDGE.
PLOT THE TWO POINTS (-2,5)&(4,-3)
NOW DRAW A VERTICLE LINE FROM Y=5 DOWN TO THE Y=-3 POINT.
DRAW A HORIZONTAL LINE FROM X=-5 LEFT TO THE X=-2 POINT.
YOU NOW HAVE A RIGHT TRIANGLE WITH SIDES EQUAL TO:
Y=5+3=8
X=4+2=6
USING THE PYTHAGOREON THEORUM A^2+B^2=C^2 WHERE A & B ARE THE TWO SIDES (X&Y) & C=THE HYPOTENUSE (OR IN THIS CASE THE DISTANCE BETWEEN THESE TWO POINTS).
6^2+8^2=C^2
36+64=C^2
C^2=100
C=SQRT100
C=10 ANSWER FOR THE HYPOTENUSE OR DISTANCE BETWEEN THESE 2 POINTS.